What is the electric potential for a sphere of radius is r

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
dear srikant 
the electric potential inside will be = Kqr / R ^3 
where k = 9 * 10 ^ 9 
the electric potential on the surface will be = Kq / r ^ 2 
HOPE IT CLEARS NOW 
REGARDS